# Multimodal Learning for Materials

### v2

@@ Abstract
Artificial intelligence is transforming computational materials science, improving the prediction of material properties, and accelerating the discovery of novel materials. Recently, publicly available material data repositories have grown rapidly. This growth encompasses not only more materials, but also a greater variety and quantity of their associated properties. Existing machine learning efforts in materials science focus primarily on single-modality tasks, i.e., relationships between materials and a single physical property, thus not taking advantage of the rich and multimodal set of material properties. Here, we introduce Multimodal Learning for Materials (MultiMat), which enables self-supervised multi-modality training of foundation models for materials. We demonstrate our framework’s potential using data from the Materials Project database on multiple axes: (i) MultiMat achieves state-of-the-art performance for challenging material property prediction tasks; (ii) MultiMat enables novel and accurate material discovery via latent space similarity, enabling screening for stable materials with desired properties; and (iii) MultiMat encodes interpretable emergent features that may provide novel scientific insights.
**footnotetext: These authors contributed equally to this work.footnotetext: {cloh, soljacic}@mit.edu
@@ I Introduction
Here, we adapt CLIP to the materials domain and also extend it to multimodal pre-training with an arbitrary number of modalities. We leverage the fact that materials databases are inherently multimodal: e.g., besides the crystal structure, the density of states (DOS) Toriyama et al. (2022); Lee et al. (2023) and charge density Dos Santos (2020) convey rich and complementary information about materials. Textual descriptions of the crystal, which can be machine-generated Rubungo et al. (2023), offer a fourth, complementary modality that is additionally computationally cheap to acquire. Motivated by these opportunities, we introduce Multimodal Learning for Materials (MultiMat), a novel framework for training a foundation model for crystalline materials that allows for the incorporation of several modalities. The basis for MultiMat is a multimodal pre-training method that connects high-dimensional material properties (i.e., modalities) in a shared latent space to produce highly effective material representations that can then be transferred to various downstream tasks. Using MultiMat, we pre-train a state-of-the-art GNN on the Materials Project Jain et al. (2013) database to demonstrate its ability to produce state-of-the-art foundation models for materials. Very recently, preliminary work has explored related ideas for molecules Takeda et al. (2023) and a structure-agnostic multi-task learning approach for crystals Prein et al. (2023).
@@ II.1 Modalities and Architecture
To illustrate the MultiMat framework, we consider four modalities for each material, all from the Materials Project database (i) the crystal structure, which we denote by $C=(\{(\mathbf{r}_{i},E_{i})\}_{i},\{\mathbf{R}_{j}\}_{j})$, where $\{(\mathbf{r}_{i},E_{i})\}_{i}$ is a set containing the coordinates $\mathbf{r}_{i}$ and chemical element $E_{i}$ of the $i$-th atom in the unit cell, and $\{\mathbf{R}_{j}\}_{j}$ is the set of unit cell lattice vectors; (ii) the DOS, $\rho(E)$, as a function of energy $E$; (iii) the charge density $n_{e}(\mathbf{r})$ as a function of position $\mathbf{r}$; and (iv) a textual description $T$ of the crystal obtained from Robocrystallographer Ganose and Jain (2019). For each material modality, we train a separate neural network encoder to learn a parameterized transformation from raw data to an embedding in a shared latent space. The $C$ encoder uses PotNet, a state-of-the-art GNN (Lin et al., 2023); the encoders of $\rho(E)$ and $n_{e}(\mathbf{r})$ are based on the Transformer Vaswani et al. (2023) and 3D-CNN architectures Xie et al. (2017). The $T$ encoder uses a frozen MatBERT Walker et al. (2021) model, a Bidirectional Encoder Representations from Transformers (BERT) textual model Devlin et al. (2019) that has been pre-trained on material science literature. A key advantage of the $T$ modality is that its data collection is relatively low cost, since Robocrystallographer can be used to generate a $T$ modality for every $C$; thus $T$ can be used to obtain a much larger pre-training dataset. Conversely, $T$ may not contain as rich information as other “high-cost” modalities like $\rho(E)$ and $n_{e}(\mathbf{r})$, which are usually obtained from ab initio simulations. Additional modality and architecture details are provided in the Methods section.
@@ II.3 Crystal Property Prediction
After the multimodal alignment stage in MultiMat, the $C$ encoder can be fine-tuned on various predictive tasks by attaching a randomly initialized linear head and fine-tuning end-to-end. We explore the tasks of predicting the bulk modulus, shear modulus, elastic tensor, and band gap corresponding to a crystal input. The mechanical property tasks use the Materials Project database Jain et al. (2013) and the band gap task uses the SNUMAT semiconductor database Kim et al. (2020). In all tasks, the fine-tuning stage has access to only relatively few labeled data points compared to the roughly 154 000154000154\,000154 000 data points used during pre-training; specifically, there are roughly $7000$ data points for the bulk modulus, shear modulus, and elastic tensor tasks and roughly 10 0001000010\,00010 000 data points for the band gap task.
@@
Figure 2 compares MultiMat using multimodal pre-training with 2 – 4range2424start_ARG 2 end_ARG – start_ARG 4 end_ARG modalities against baselines without multimodal pre-training. The two baselines are CGCNN (Xie and Grossman, 2018) (a pioneering method in the use of GNNs for crystal property prediction) and PotNet (Lin et al., 2023), the current state-of-the-art method for crystal property prediction using GNNs. Note that MultiMat also uses the architecture of PotNet for the $C$ encoder. For the two- and three-modality cases in Fig. 2, the shown results are averages of experiments over all possible two- or three-modality combinations from the set $\{C,\rho(E),n_{e}(\mathbf{r}),T\}$ with $C$ always chosen. For example, for two modalities, results are the average MAE over the combinations $\{C,\rho(E)\}$, $\{C,n_{e}(\mathbf{r})\}$, and $\{C,T\}$ (results of individual experiments are shown in the Supplementary Information). MultiMat significantly improves predictive performance over the baselines, with up to  ${\sim}\,10\%$ reduction in MAE over the best baseline (PotNet).
@@
We observe that including three or more modalities during pre-training improves performance over just two modalities (see Fig. 2). On the other hand, we found no significant gains in using MultiMat with $M=4$ modalities over $M=3$ modalities. Speculatively, this might reflect that: (i) there is only marginal additional information from the fourth modality or (ii) the current implementation is not able to take advantage of the additional information due to model capacity limitations (to ensure fair comparison, we use a fixed architecture across all experiments and do not increase the neural network capacity to match the corresponding increased complexity).
@@ II.4 Material Discovery via Latent Space Similarity
The alignment between modalities MultiMat optimizes for ensures that a close match between $C$ and $\rho(E)$ embeddings in the multimodal space signifies similarity in the physical space between the candidate material corresponding to the $C$ embeddings and the material corresponding to the target $\rho(E)$. This proposed material discovery approach leverages the extensive scale of $C$ databases, which typically exceeds the number of entries for other modalities by at least an order of magnitude, and thus allows one to identify existing materials that would have a $\rho(E)$ very similar to the target, had it been computed. This constitutes an accelerated form of material design, which only uses inference through the neural network encoders followed by a nearest neighbor search to find materials likely to exhibit certain desired properties. This material discovery approach is enabled by the alignment between encoders that MultiMat optimizes. In contrast “forward-only” approaches to material design are based on an encoder-decoder structure (e.g., where $C$ is first encoded to a latent space and then decoded to predict $\rho(E)$). A potential benefit of our latent-based similarity approach lies in the fact that searching for candidates in a low-dimensional latent space compared to the physical space is likely easier for high-dimensional properties such as $\rho(E)$. A related work focusing on $\rho(E)$ was introduced in Ref. 64; however, it differs from ours by only working for binary composition materials and focusing on material design through chemical composition for a fixed atomic structure, thus neglecting the structural information of materials. Note that while our results focus on $\rho(E)$, the approach is applicable to other modalities, provided that the respective encoders are trained with MultiMat.
@@ IV.3 Material Discovery via Latent Space Similarity and Interpretability of Embeddings
Note that this metric, despite its relative character, may still exhibit large values (e.g., exceeding unity), even for slight misalignment of the resonance energies because the DOS frequently is a sharply peaked quantity.

### v4

# This article has been accepted for publication in Newton (Cell Press)

@@ Abstract
Artificial intelligence is transforming computational materials science, improving the prediction of material properties, and accelerating the discovery of novel materials. Recently, publicly available material data repositories have grown rapidly. This growth encompasses not only more materials but also a greater variety and quantity of their associated properties. Existing machine learning efforts in materials science focus primarily on single-modality tasks, i.e., relationships between materials and a single physical property, thus not taking advantage of the rich and multimodal set of material properties. Here, we introduce Multimodal Learning for Materials (MultiMat), which enables self-supervised multi-modality training of foundation models for materials. We demonstrate our framework’s potential using data from the Materials Project database on multiple axes: (i) MultiMat achieves state-of-the-art performance for challenging material property prediction tasks; (ii) MultiMat enables novel and accurate material discovery via latent space similarity, enabling screening for stable materials with desired properties; and (iii) MultiMat encodes interpretable emergent features that may provide novel scientific insights.
**footnotetext: These authors contributed equally to this work.footnotetext: {vmoro, soljacic}@mit.edu
@@ I Introduction
Here, we adapt CLIP to the materials domain and also extend it to multimodal pre-training with an arbitrary number of modalities. We leverage the fact that materials databases are inherently multimodal: e.g., besides the crystal structure, the density of states (DOS) Toriyama et al. (2022); Lee et al. (2023) and charge density Dos Santos (2020) convey rich information about materials. Textual descriptions of the crystal, which can be machine-generated Rubungo et al. (2023), offer a fourth modality that is additionally computationally cheap to acquire. It is important to point out that the above-mentioned material modalities are not information-independent, since they can be computed from the crystal structure. The same holds true for image-caption pairs that were used in CLIP. Therefore, the point of contrastive multimodal pre-training is not to leverage modalities with independent information but rather to learn better representations by integrating different perspectives of the same underlying data Wang and Isola (2020); Chen et al. (2020); van den Oord et al. (2019); Daunhawer et al. (2023). Motivated by these opportunities, we introduce Multimodal Learning for Materials (MultiMat), a novel framework for training a foundation model for crystalline materials that allows for the incorporation of several modalities. The basis for MultiMat is a multimodal pre-training method that connects high-dimensional material properties (i.e., modalities) in a shared latent space to produce highly effective material representations that can then be transferred to various downstream tasks. Using MultiMat, we pre-train a state-of-the-art GNN on the Materials Project Jain et al. (2013) database to demonstrate its ability to produce state-of-the-art foundation models for materials. Very recently, preliminary work has explored related ideas for molecules Takeda et al. (2023) and a structure-agnostic multi-task learning approach for crystals Prein et al. (2023).
@@ II.1 Modalities and Architecture
To illustrate the MultiMat framework, we consider four modalities for each material, all from the Materials Project database (i) the crystal structure, which we denote by $C=(\{(\mathbf{r}_{i},E_{i})\}_{i},\{\mathbf{R}_{j}\}_{j})$, where $\{(\mathbf{r}_{i},E_{i})\}_{i}$ is a set containing the coordinates $\mathbf{r}_{i}$ and chemical element $E_{i}$ of the $i$-th atom in the unit cell, and $\{\mathbf{R}_{j}\}_{j}$ is the set of unit cell lattice vectors; (ii) the DOS, $\rho(E)$, as a function of energy $E$; (iii) the charge density $n_{e}(\mathbf{r})$ as a function of position $\mathbf{r}$; and (iv) a textual description $T$ of the crystal obtained from Robocrystallographer Ganose and Jain (2019). For each material modality, we train a separate neural network encoder to learn a parameterized transformation from raw data to an embedding in a shared latent space. The $C$ encoder uses PotNet, a state-of-the-art GNN (Lin et al., 2023); the encoders of $\rho(E)$ and $n_{e}(\mathbf{r})$ are based on the Transformer Vaswani et al. (2023) and 3D-CNN architectures Xie et al. (2017). The $T$ encoder uses a frozen MatBERT Walker et al. (2021) model, a Bidirectional Encoder Representations from Transformers (BERT) textual model Devlin et al. (2019) that has been pre-trained on material science literature. A key advantage of the $T$ modality is that its data collection is relatively low cost since Robocrystallographer can be used to generate a $T$ modality for every $C$; thus $T$ can be used to obtain a much larger pre-training dataset. Conversely, $T$ may not contain as rich information as other “high-cost” modalities like $\rho(E)$ and $n_{e}(\mathbf{r})$, which are usually obtained from ab initio simulations. Additional modality and architecture details are provided in the Methods section.
@@ II.3 Crystal Property Prediction
After the multimodal alignment stage in MultiMat, the $C$ encoder can be fine-tuned on various predictive tasks by attaching a randomly initialized linear head and fine-tuning end-to-end. We explore the tasks of predicting the bulk modulus, shear modulus, elastic tensor, and band gap corresponding to a crystal input. The mechanical property tasks use the Materials Project database Jain et al. (2013) and the band gap task uses the SNUMAT semiconductor database Kim et al. (2020). These tasks were chosen because they have relatively few labeled data points compared to the number of data points used during pre-training. In particular, roughly 154 000154000154\,000154 000 data points are used during pre-training compared to roughly $7000$ data points for the bulk modulus, shear modulus, and elastic tensor tasks and roughly 10 0001000010\,00010 000 data points for the band gap task. Note that for the crystal property prediction tasks, only the crystal structure is used (and not any of the other modalities used for multimodal pre-training).
@@
Figure 2 compares MultiMat using multimodal pre-training with 2 – 4range2424start_ARG 2 end_ARG – start_ARG 4 end_ARG modalities against baselines without multimodal pre-training. The two baselines are CGCNN (Xie and Grossman, 2018), the first method using GNNs for crystal property prediction, and PotNet (Lin et al., 2023), the current state-of-the-art method for crystal property prediction using GNNs. Note that MultiMat also uses the architecture of PotNet for the $C$ encoder. For the two- and three-modality cases in Fig. 2, the shown results are averages of experiments over all possible two- or three-modality combinations from the set $\{C,\rho(E),n_{e}(\mathbf{r}),T\}$ with $C$ always chosen. For example, for two modalities, results are the average MAE over the combinations $\{C,\rho(E)\}$, $\{C,n_{e}(\mathbf{r})\}$, and $\{C,T\}$ (results of individual experiments are shown in the Supplementary Information). MultiMat pre-training significantly improves predictive performance compared to the baselines that do not make use of any pre-training. In particular, MultiMat reduces the MAE by up to  ${\sim}\,10\%$ compared to PotNet, which is the current state-of-the-art and the $C$ architecture used in MultiMat. This performance improvement is comparable to the improvement when going from CGCNN to PotNet, methods that are separated by five years that represent the first method and state-of-the-art method for crystal property prediction respectively. Moreover, note that MultiMat is a pre-training method that can be used on top of any existing or future crystal encoder to substantially improve its performance. Thus, the primary focus is on the performance difference between PotNet and MultiMat, with the CGCNN baseline serving to contextualize the overall performance advancements.
@@
We observe that including three or more modalities during pre-training marginally improves performance over just two modalities (see Fig. 2). On the other hand, we found no significant gains in using MultiMat with $M=4$ modalities over $M=3$ modalities. Speculatively, this might reflect that: (i) the fourth modality offers only a marginally different perspective on the material compared to the other three modalities or (ii) the current implementation is not able to take advantage of the additional modality due to model capacity limitations (to ensure fair comparison, we use a fixed architecture across all experiments and do not increase the neural network capacity to match the corresponding increased complexity).
@@ II.4 Material Discovery via Latent Space Similarity
The alignment between modalities MultiMat optimizes for ensures that a close match between $C$ and $\rho(E)$ embeddings in the multimodal space signifies similarity in the physical space between the candidate material corresponding to the $C$ embeddings and the material corresponding to the target $\rho(E)$. This proposed material discovery approach leverages the extensive scale of $C$ databases, which typically exceeds the number of entries for other modalities by at least an order of magnitude, and thus allows one to identify existing materials that would have a $\rho(E)$ very similar to the target, had it been computed. This constitutes an accelerated form of material design, which only uses inference through the neural network encoders followed by a nearest neighbor search to find materials likely to exhibit certain desired properties. This material discovery approach is enabled by the alignment between encoders that MultiMat optimizes. In contrast “forward-only” approaches to material design are based on an encoder-decoder structure (e.g., where $C$ is first encoded to a latent space and then decoded to predict $\rho(E)$). A potential benefit of our latent-based similarity approach lies in the fact that searching for candidates in a low-dimensional latent space compared to the physical space is likely easier for high-dimensional properties such as $\rho(E)$. A related work focusing on $\rho(E)$ was introduced in Ref. 68; however, it differs from ours by only working for binary composition materials and focusing on material design through chemical composition for a fixed atomic structure, thus neglecting the structural information of materials. Note that while our results focus on $\rho(E)$, the approach is applicable to other modalities, provided that the respective encoders are trained with MultiMat.
@@ IV.3 Material Discovery via Latent Space Similarity and Interpretability of Embeddings
Note that this metric, despite its relative character, may still exhibit large values (e.g., exceeding unity), even for a slight misalignment of the resonance energies because the DOS frequently is a sharply peaked quantity.
@@ IV.5 Implementation Details and Settings for Training and Evaluation
IV.6 Data availability
This article uses data that are all open-sourced and publicly available. Specifically, the data is downloaded from the Materials Project (https://next-gen.materialsproject.org/) and SNUMAT (https://www.snumat.com/) databases.
IV.7 Code availability
MultiMat was developed using the PyTorch framework. All source codes used for training and for generating all the results are publicly available at https://github.com/vmoro1/multimat.